OpenStudy (anonymous):

A mathematics journal has accepted 14 articles for publication. However, due to budgetary restraints only 9 articles can be published this month. How many ways can the journal editor assemble 9 of the 14 articles for publication? A. 726,485,760 B. 14 C. 2,002 D. 126

OpenStudy (anonymous):

you are being asked for the number of ways to choose 9 out of 14 this is written as \(\dbinom{14}{9}\) read " fourteen choose 9" and compute via \[\frac{14\times 13\times 11\times 10\times 9}{5\times 4\times 3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

typo \[\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

excellent answer..its absolutely right..

OpenStudy (anonymous):

So the answer is C

OpenStudy (anonymous):

yes